The force between two identical point charges in F.  What will be the force if each charge is increased by 2 % and the distance between them is decreased by 3% ? 

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
let r be the distance b/w the charges.
Then,
F = kq^2/r^2........(1)
After the changes in cahrges and the distance the new Force becomes:
F’= kq^2(1.02)^2/(0.98r)^2...............(2)
Divide both the eqns. we get,
F/F’= (0.98/1.02)^2
F’= (1.02/0.98)^2*F => 1.083F This would be the new Force.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)